Two vehicles A (mass 520kg) and B (1.7 tons) collide at an intersection. Before the collision, vehicle A traveled at a velocity of 120km/h in an eastern direction and vehicle B traveled with a velocity of 85km/h in a northern direction. The collision is in elastic. Calculate the momentum of Vehicle A before the collision (magnitude only
1
Expert's answer
2020-09-14T10:20:03-0400

"p_A=m_Av_A=520\cdot 33.33=17333(\frac{kg\cdot m}{s})"